An electric dipole consist of two xharges of 0.1 macro coulomb separated by a distanceof 2 cm.the dipole is placed in a external field of a 10^5N/C.what maximum torque does the field exert on the dipole.

Grade:12

1 Answers

Ojas Bisariya
7 years ago
Maximum torque = PEsin90°Now, P = total charge × effective length =0.1×10^-6 × 0.02 P = 2× 10^-9 CmAnd E = 10^5 N/CTorque = 2×10^-9×10^5×1 (•.• sin90= 1) Torque = 2×10^-4 Nm